OpenStudy (anonymous):

Can anybody please help. It is hard. Luisa mows lawns during the summer. She charges $15 if she cuts the grass but charges $5 more if she also trims the grass. Last week she trimmed 5 more yards than she cut. If she made $415 last week, how many yards did she trim? Please explain step by step.Thanks!

OpenStudy (anonymous):

Thanks for viewing please reply.

OpenStudy (anonymous):

if x = number of yards just cut and x + 5 = number of yards cut AND trimmed then 15x + 20(x+5) = 415. from this you can set it up to 15x + 20x+100 = 415. now just simplify this equation x = 315/35 x = 9 (this is just cut) this means that the yards trimmed and cut is x+5 = 9+5 = 14